what is the probability of having exactly 3 girls out of five children?

Answer by stanbon(75887)

what is the probability of having exactly 3 girls out of five children?
---------------------------
P(x = 3) = 5C3*(1/2)^3*(1/2)^2 = 10(1/32) = 10/32 = 5/16
=========================================================
Cheers,
Stan H.